Calculate the value of Kp for the equation


C(s) + CO2(g) <--> 2CO(g) Kp = ?


Given that at a certain temperature


C(s) + 2H2O(g) <--> CO2(g) + 2H2(g) Kp1 = 3.97


H2(g) + CO2(g) <--> H20(g) + CO(g) Kp2 = 0.623
Answer #1
Concepts and reason

An equilibrium constant for a chemical reaction is a certain value of the reaction at equilibrium state, which depends on the temperature at equilibrium. The equilibrium constant for an equilibrium reaction can be expressed in two ways: one is in terms of concentration, the other is in terms of partial pressures.

Fundamentals

The equilibrium constant in terms of partial pressure is represented as. The equilibrium constant can be represented for both homogeneous and heterogeneous equilibria.

Homogeneous equilibria:

The reactants and products are in the same phase.

Example:

A(g) + B(g) =
C(g)+D(8)

Heterogeneous equilibria:

The reactants and products are in different phases.

A(8) + B(g) =
C(g)+D(S)

The relationship between two equilibrium constants can be:

K,=K (RT)An

C(s) +24,0(9)
H2(g)+C02(9)=
CO2(g)+2H,(9) —>1 Kp,=3.97
H2O(g)+CO(g) →→2 Kpz=0.623
equation 2 is multiplied by 2:
C(s) +22,0(9

According to the given reactions:
Kp = Kp, x(Kp2)
= 3.97*(0.623)
Kp = 1.540

Ans:

C(s) + CO2(g)
200(g)
Kp = 1.540